i need your help
16/9=42/x

Found 2 solutions by stanbon, katz:
Answer by stanbon(75887)

16/9=42/x
-------
both sides to get:
x/42 = 9/16
x=42(9/16)
x=21(9/8)
x=23.625 or 189/8
Cheers,
Stan H.

Answer by katz(9)

16/9=42/x
x=189/8
16/9=42/x
16/9=42/189/8
16/9=2^4*3*7/3^3*7
16/9=16/9
x=189/8